R1 = 6 ohm

R2 = 5 ohm

R3 = 4 ohm

R4 = 4 ohm

R5 = 8 ohm

with 10V

Solve for the voltage across resistor #5 (R5)?


1
Expert's answer
2021-03-15T20:58:31-0400

Assume the circuit is as shown below





"R_2" and "R_3" are in series


"R_{2}+R_3=5+4=9\Omega"


"R_2,R_3and R_4" combined gives

"1/R=1/9+1/4=13/36, R=36/13=2.769\Omega"

Effective resistance, "R_E=6+2.769+8=16.769\Omega"

given "v=10v"

"I=V\/R_E=10\/16.769=0.596A"

"V_{R5}=IR=0.596\times 8=4.768v"